OpenStudy (anonymous):

What is the solution set for |-2x + 8| = 12

OpenStudy (anonymous):

answers... x = -2 and x = 10 x = -2 and x = -10 x = 2 and x = 10 x = 2 and x = -10

OpenStudy (ajprincess):

|a|=b a=b or a=-b Similarly |-2x + 8| = 12 -2x+8=12 or -2x+8=-12 Can u solve nw @jgoodman?

OpenStudy (anonymous):

@ajprincess yes, thank you